OpenStudy (anonymous):

if 3x^2-2x+7=0, then (x-1/3)^2=?

OpenStudy (anonymous):

3x^2-2x+7=0 x^2 - (2/3)x + (7/3) = 0 (x -1/3)^2 - 1/9 + 7/3 = 0 (x - 1/3)^2 = - 20/9

OpenStudy (anonymous):

i see that you divided by 3 in the second equation but how did you derived at the third equation?

OpenStudy (anonymous):

That is called "completing the square" x^2 - (2/3)x = (x -1/3)^2 - 1/9 Notice that when you expand (x -1/3)^2, you will get x^2 - (2/3)x + 1/9. Is this help?